OpenStudy (anonymous):

9 + (8 + 1) = (9 + 8) + 1

OpenStudy (akashdeepdeb):

This is known as the associative property of addition in arithmetic. Where a + (b+c) = (a+b) + c :D

OpenStudy (hang254):

9 + (8 +1) = (9+8) +1 9 + (9) = (17) +1 18=18

OpenStudy (hang254):

jut remember PEMDAS parenthesis, exponents, multiply, divide, add, subtract.